OpenStudy (anonymous):

Four more than twice a number is seven more than the number. What is the number?

OpenStudy (anonymous):

hi:DD

OpenStudy (a1234):

2n + 4 = n + 7 Subtract n from both sides: n + 4 = 7 Subtract 4 from both sides: n = 3 Check to see if it works.

OpenStudy (anonymous):

it does, thank you!!! (: